OpenStudy (anonymous):

Suppose that 74% of all students at a given school have their own cell phone. If five students are surveyed at random a) What is the probability at least two of them have their own cell phone? b) What is the expected number of students with their own cell phone?

OpenStudy (anonymous):

well for part a P(a) = cos(74%) + sin(74%) P(a | CellPhoneAvailable) = 1 - P(a) = 0.982 so final answer is .982

OpenStudy (anonymous):

part b just multiply .74 * 5 = 3.7

OpenStudy (anonymous):

ok thnx